Earned Value Management (EVM) Following is a cost/schedule status report (C/SSR), a document that is generated...

  1. Earned Value Management (EVM)

Following is a cost/schedule status report (C/SSR), a document that is generated each month. It demonstrates the cost and schedule status of a project at a given moment in time. Answer the questions associated with the accompanying C/SSR after reading the appropriate sections on earned value in Managing Projects in Organizations, The New Project Management, and A Guide to the Project Management Body of Knowledge (PMBOK).

Task

Budget

Task begun

Task ended

Actual Cost

A

30,000

30,000

B

40,000

43,000

C

20,000

22,000

D

30,000

18,000

E

30,000

F

40,000

44000

G

20,000

21,000

Total

210,000

178,000

This C/SSR shows the cost and schedule status of our project as of the end of last month. The numbers in the “Budget” column represent how much money was budgeted to be spent on each of the tasks at the end of reporting period. Applying earned value management (EVM) principles, answer the following six questions:

  1. What is the schedule variance (SV) for the project as of the end of last month?
  2. What is the cost variance (CV) for the project as of the end of last month?
  3. What is the value of cost performance index (CPI)?
  4. What is the value of the schedule performance index (SPI)?
  5. If the total budget for the project as of its completion date is $600,000, what is the value of estimate at complete (EAC)?
  6. In one paragraph, summarize the cost and schedule status of this project using the earned value
Answer #1

AC = actual cost,EV = Earned value,PV = Planned value

a)

SV = EV-PV

EV = (4/7)*210000 = 120000

PV = 210000

SV = EV-PV = 120000-210000 = -90000

b)

CV = EV-AC = 120000-178000 = -58000

c)

CPI = EV/AC = 120000/178000 = 0.674157303

d)
SPI = EV/PV = 120000/210000 = 0.571428571

e)
EAC = budget at completion + AC -EV

=600000 + 178000 – 120000 = 658000

f)

CPI < 1 suggests that project is over budget which means project may run out of money before completion. SPI < 1 suggests the project is behind schedule

CPI < 1